# Brainformer: Modeling MRI Brain Functions to Machine Vision

### v1

@@ 3.3.2 Brain fMRI Guidance Loss
Algorithm 1 Mutli-scale fMRI Transformer
1:  Input: The feature $\mathbf{r}^{k}$ of $k^{th}$ fMRI signals , window size $w$, stride: $s$, number of level $h$.
2:  Output: The fMRI features $\bar{\mathbf{q}}^{k}$
3:  $\mathbf{x}\leftarrow\mathbf{r}^{k}$
4:  for $j=1,\cdots,h$ do
[…]
6:     $L\leftarrow|x|$
7:     for $i=0,\cdots\frac{L}{s}$ do
8:        $\mathbf{q}_{i}^{k}\leftarrow\mathbf{x}\left[i*s:i*s+w\right]$triangleright Getting subsequences
9:        $\bar{\mathbf{q}}_{i}^{k}\leftarrow\texttt{TransBlock}^{k}_{j}(\mathbf{q}_{i}^{k})$
10:        $\mathbf{t}\leftarrow\left[\mathbf{t}\quad\bar{\mathbf{q}}_{i}^{k}\right]$triangleright Gather features of the subsequence
11:     end for
12:     $\mathbf{x}\leftarrow\mathbf{t}$
13:  end for
14:  $\bar{\mathbf{q}}^{k}\leftarrow\texttt{TransBlock}(\mathbf{x})$
15:  return $\bar{\mathbf{q}}^{k}$
@@
$\small\begin{split}\mathcal{L}_{bfg}&=-\frac{1}{N}\sum_{i}^{N}\sum_{k}^{N_{r}}\log\frac{\exp(\bar{\mathbf{p}}^{k}_{i}\otimes\bar{\mathbf{q}}^{k}_{i})}{\sum_{g}^{N_{r}}\exp(\bar{\mathbf{p}}^{k}_{i}\otimes\bar{\mathbf{q}}^{g}_{i})}\end{split}$ (7)
@@ 4.1 Datasets
To pretrain Brainformer, we leverage the Natural Scenes Dataset (NSD) [1], a comprehensive compilation of responses from eight subjects obtained through high-quality 7T fMRI scans. Each subject was exposed to approximately 73,000 natural scenes, forming the basis for constructing visual brain encoding models. The fMRI responses of each image encompass the left hemisphere (LH) and right hemisphere (RH), with 19,004 and 20,544 voxels, respectively.
@@ 4.2 Brainformer Training
We use data from seven subjects in NSD for training and leave one for testing. The images are resized to 224 $\times$ 224 before feeding into the vision model. We do not use any augmentations for the image because the fMRI signals depend on the input image. Any change in vision stimuli will affect human recognition. For the fMRI signals, we divide them into six regions of interest following [16] and feed them into Brainformer simultaneously. Experimentally, we select Conv1D with kernel size and stride are 32 and 16, respectively. The Multi-scale fMRI Transformer is designed with a window size as $w=64$, stride as $s=32$, and number of levels as $h=2$. Feature dimensions of image and fMRI signals are projected into $d_{r}=768$ space. For the vision model, we select Swin-S [34] and ConvNext-S [35] as the image backbone. Brainformer is easily implemented in the Pytorch framework and trained on A100 GPUs. The initial learning rate is 0.0001 and decreases gradually following the ConsineLinear policy [36]. A batch size of 64 per GPU is employed. Optimization is performed using AdamW [37] with 100 epochs, with the entire training process concluding within a few hours. The pre-trained models are used for further downstream tasks, including COCO object detection [32] and ADE20K semantic segmentation [69].
@@ 4.3 Object Detection on COCO
  (a) Object Detection
[…]
Swin-S Random init 41.3 63.4 45.5
Swin-S CLIP [48] 41.9 64.0 46.0
Swin-S Brainformer 43.6 65.8 47.4
ConvNext-S Random init 42.6 65.5 47.0
ConvNext-S CLIP [48] 42.8 66.1 48.3
ConvNext-S Brainformer 45.1 68.2 50.0
(b) Semantic Segmentation
[…]
Swin-S Random init 38.4 60.8 41.3
Swin-S CLIP[48] 39.2 61.3 41.7
Swin-S Brainformer 43.1 63.1 43.3
ConvNext-S Random init 39.8 61.7 42.9
ConvNext-S CLIP [48] 40.8 62.2 43.5
ConvNext-S Brainformer 44.0 64.0 45.3
@@ 4.4 Semantic Segmentation on ADE20K
Swin-S Random init 38.37 40.41
Swin-S CLIP [48] 40.29 41.25
Swin-S Brainformer 41.77 44.63
ConvNext-S Random init 39.22 54.21
ConvNext-S CLIP [48] 41.27 55.70
ConvNext-S Brainformer 42.92 57.43
@@ 5 Ablation Studies
Figure 6: Visual attention with respect to fMRI signals. The first row is the input images. The second row is the corresponding attention maps of the vision model training with CLIP. The third row is the results of training with Brainformer without Brain fMRI Guidance Loss. The last row is the results with Brainformer and guidance of Brain fMRI Guidance Loss. Best view in color.
Table 3: Performance of Brainformer on various settings.

### v2

@@ 3.3.2 Brain fMRI Guidance Loss
$\begin{split}\mathcal{L}_{bfg}&=-\frac{1}{N}\sum_{i}^{N}\sum_{k}^{N_{r}}\log\frac{\exp(\bar{\mathbf{p}}^{k}_{i}\otimes\bar{\mathbf{q}}^{k}_{i})}{\sum_{g}^{N_{r}}\exp(\bar{\mathbf{p}}^{k}_{i}\otimes\bar{\mathbf{q}}^{g}_{i})}\end{split}$ (6)
@@ 4.1 Datasets
To pretrain Brainformer, we leverage the Natural Scenes Dataset (NSD) [1], a comprehensive compilation of responses from eight subjects obtained through high-quality 7T fMRI scans. Each subject was exposed to approximately 73,000 natural scenes, forming the basis for constructing visual brain encoding models.
@@ 4.2 Brainformer Training
We use data from seven subjects in NSD for training and leave one for testing. The images are resized to 224 $\times$ 224 before feeding into the vision model. We do not use any augmentations for the image because the fMRI signals depend on the input image. Any change in vision stimuli will affect human recognition. For the fMRI signals, we divide them into six regions of interest following [20] and feed them into Brainformer simultaneously. Experimentally, we select Conv1D with kernel size and stride are 32 and 16, respectively.
The Multi-scale fMRI Transformer is designed with a window size as $w=64$, stride as $s=32$, and number of levels as $h=2$. Feature dimensions of image and fMRI signals are projected into $d_{r}=768$ space. For the vision model, we select Swin-S [45] and ConvNext-S [46] as the image backbone. Brainformer is easily implemented in the Pytorch framework and trained on 8 $\times$ A100 GPUs. The initial learning rate is 0.0001 and decreases gradually following the ConsineLinear policy [47]. A batch size of 64 per GPU is employed. Optimization is performed using AdamW [48] with 100 epochs, with the entire training process concluding within two hours. The pre-trained models are used for further downstream tasks, including COCO object detection [42] and ADE20K semantic segmentation [98].
                           (a) Object Detection
[…]
Swin-S Random init 41.3($\pm$0.2) 63.4($\pm$0.3) 45.5($\pm$0.4)
Swin-S CLIP [72] 41.9($\pm$0.1) 64.0($\pm$0.4) 46.0($\pm$0.3)
Swin-S Brainformer 43.6($\pm$0.2) 65.8($\pm$0.5) 47.4($\pm$0.3)
ConvNext-S Random init 42.6($\pm$0.3) 65.5($\pm$0.2) 47.0($\pm$0.2)
ConvNext-S CLIP [72] 42.8($\pm$0.1) 66.1($\pm$0.2) 48.3($\pm$0.3)
ConvNext-S Brainformer 45.1($\pm$0.1) 68.2($\pm$0.2) 50.0($\pm$0.2)
(b) Semantic Segmentation
[…]
Swin-S Random init 38.4($\pm$0.2) 60.8($\pm$0.3) 41.3($\pm$0.2)
Swin-S CLIP[72] 39.2($\pm$0.3) 61.3($\pm$0.3) 41.7($\pm$0.2)
Swin-S Brainformer 43.1($\pm$0.2) 63.1($\pm$0.5) 43.3($\pm$0.2)
ConvNext-S Random init 39.8($\pm$0.4) 61.7($\pm$0.1) 42.9($\pm$0.3)
ConvNext-S CLIP [72] 40.8($\pm$0.4) 62.2($\pm$0.2) 43.5($\pm$0.1)
ConvNext-S Brainformer 44.0($\pm$0.2) 64.0($\pm$0.5) 45.3($\pm$0.4)
@@ 4.4 Semantic Segmentation on ADE20K
Swin-S Random init 38.37($\pm$0.2) 40.41($\pm$0.4)
Swin-S CLIP [72] 40.29($\pm$0.3) 41.25($\pm$0.3)
Swin-S Brainformer 41.77($\pm$0.3) 44.63($\pm$0.2)
ConvNext-S Random init 39.22($\pm$0.3) 54.21($\pm$0.3)
ConvNext-S CLIP [72] 41.27($\pm$0.4) 55.70($\pm$0.2)
ConvNext-S Brainformer 42.92($\pm$0.3) 57.43($\pm$0.4)
@@ 5 Ablation Studies
Table 3: Performance on various settings.
@@
Figure 6: Visual attention with respect to fMRI signals. The first row is the input images. The second row is the corresponding attention maps of the vision model training with CLIP. The third row is the results of training with Brainformer without Brain fMRI Guidance Loss. The last row is the results with Brainformer and guidance of Brain fMRI Guidance Loss. The warmer colors, the higher attention. Best view in color.